OpenStudy (raffle_snaffle):

What mass of silver chloride can be produced from 1.20 L of a 0.259 M solution of silver nitrate? answer for first part is 4.45g The reaction described in Part A required 3.00 L of potassium chloride. What is the concentration of this potassium chloride solution?

OpenStudy (raffle_snaffle):

@Preetha

OpenStudy (raffle_snaffle):

not 4.45 g it's suppose to be 44.5g

OpenStudy (kkutie7):

usually when concentration is being asked for you need to convert to Moles/L

OpenStudy (raffle_snaffle):

^That was of no help....

OpenStudy (anonymous):

\[KCl+AgNO _{3}\rightarrow AgCl+KNO _{3}\] Write equation first.

OpenStudy (anonymous):

To find the ratio of moles or volume. First find number of moles of the compound with more data that facilitates the calculation. It will be of Silver Nitrate.\[mol=\frac{ Concentration*Volume (cm^3) }{ 1000 }=\frac{ 0.259*1200 }{ 1000 }=0.3108mol\] According to the equation, the ratio of moles of AgNO3 to KCl is 1:1 It means KCl should have the same number of moles of AgNO3, that is 0.3108mol. Re-use the same formula to make concentration the subject. That will be:\[concentration=\frac{ 1000mol }{ Volume(cm^3) }=\frac{ 1000*0.3108 }{ 3000 }=0.1036M\]